OpenStudy (anonymous):

the volume of 100 drops of a liquid is 1 fluid ounce. What is the volume of 10 drops?

OpenStudy (texaschic101):

100/1 = 10/x --- (100 drops to 1 fl oz = 10 drops to x fl oz) cross multiply (100)(x) = (1)(10) 100x = 10 x = 10/100 x = 1/10 fl oz 10 drops = 1/10 fl oz

OpenStudy (anonymous):

Thank you!